OpenStudy (angiepangie0726):

Solve 6x3 − x2 − 27x − 20 = 0 by factoring.

OpenStudy (angiepangie0726):

@Ms-Brains Sorry about the last post, it was really hard.

MsBrains (ms-brains):

Is this Multiple choice?

OpenStudy (angiepangie0726):

Sadly no :(

MsBrains (ms-brains):

Maybe this can help then http://www.wolframalpha.com/input/?i=6x3+%E2%88%92+x2+%E2%88%92+27x+%E2%88%92+20+%3D+0

OpenStudy (angiepangie0726):

I need step by step and it didnt give it to me that way:/

OpenStudy (angiepangie0726):

@phi

OpenStudy (angiepangie0726):

Do you have any idea what it may be? @Ms-Brains

OpenStudy (angiepangie0726):

@Stacey

OpenStudy (angiepangie0726):

@563blackghost

OpenStudy (angiepangie0726):

@ILovePuppiesLol

OpenStudy (angiepangie0726):

@agent0smith @sweetburger

OpenStudy (angiepangie0726):

Can anyone help?

OpenStudy (stacey):

This does factor into three binomials. I am looking at how to explain the factoring. Have you learned synthetic division?

alones (alones):

does it looks liek this \(6x^{3}\) − x2 − 27x − 20 = 0

OpenStudy (anonymous):

\[6x^3-x^2-27x-20=0\] put x=-1 -6-1+27-20=-27+27=0 hence x=-1 is its zero divide by -1